OpenStudy (anonymous):

Solve the system algebraically. 4x - 5y = 3 x = 3 - 1/2y What is the value of x? 9/7 24/14 33/14

OpenStudy (anonymous):

huh?

OpenStudy (unklerhaukus):

\[4x - 5y = 3\qquad \qquad x = 3 - 1/2y\] \[4(3 - 1/2y) - 5y = 3\]\[12-4/2y-5y=3\]\[9-7y=0\]\[9=7y\]\[\frac97=y\] \[x = 3 - 1/2\left(\frac97\right)\]\[x=3-9/7\]\[=(21-9)/7\]\[x=12/7\]

OpenStudy (unklerhaukus):

b